OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 3x2 – 24x + 24 = 0

OpenStudy (anonymous):

3x2 – 24x + 24 = 0 3(x^2 -8x +8)=0 (x^2 -2(4)x +16-16+8)=0 (x-4)^2-8=0 or (x-4)^2=8 --> (x+(-4))^2=8 here c=-4 and a=8

OpenStudy (anonymous):

Thank you!